Two years ago, a political party ASR received 9:8% of the votes in an election. To study the current
political preferences, a statistical research institute plans to organise a poll by the end of the present
year. In this study, n voters will be interviewed about the political party they prefer.
Below, p, denotes the proportion of voters that would vote ASR if the elections were held now.
Furthermore, b p denotes the (random) sample proportion of the ASR voters.
1
Expert's answer
2020-03-30T11:57:52-0400

We need to construct the "90\%" confidence interval for the population proportion. We have been provided with the following information about the sample proportion:

"\begin{matrix}
 Sample\ proportion & \hat{p}=0.098 \\
 Sample\ Size & N=500
\end{matrix}"

The critical value for "\alpha=0.1" is "z_c=z_{1-\alpha/2}=1.645". The corresponding confidence interval is computed as shown below:


"CI(Proportion)=""=\big(\hat{p}-z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}},\hat{p}+z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}} \big)"

"=(\hat{0.098}-1.645\sqrt{{\hat{0.098}(1-\hat{0.098}) \over 500}},\hat{0.098}+1.645\sqrt{{\hat{0.098}(1-\hat{0.098}) \over 500}})="

"=(0.076,0.120)"

 How large should the sample size be to Obtain an interval width about 0:02?

An interval width is


"width=\hat{p}+z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}}-(\hat{p}-z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}})="

"=2z_c\sqrt{{\hat{p}(1-\hat{p}) \over N}}"

"2(1.645)\sqrt{{\hat{0.098}(1-\hat{0.098}) \over N}}=0.02"

"N={ 0.098(1-0.098)\over \big(\dfrac{0.02}{2(1.645)}\big)^2}\approx2392"